OpenStudy (anonymous):

Solve A = 2πr + 3t for t

OpenStudy (goten77):

\[A= 2\pi r + 3t \] then becomes \[A-2\pi r = 3t\] then divide by 3 and you have \[\left( A-2\pi r \right) / 3\] = t

OpenStudy (anonymous):

Thank you very much my friend is doing hw and she is asking me for the answers and Idk all of them